DVD Now is a direct marketer of popular movies. Following is information about its revenue and cost structure:

Selling Price $13 per dvd

Variable costs:

Production (manufacturing costs) $3.00 per dvd

selling and administration (non manuf. cost). $1.00 per dvd

fixed costs:

Production (manufacturing costs) $1,000,000 per year

Selling and administration ( non-manu. costs). $3,000,000 per year

In which range dos the break-even point fall?

a. between 300,000 and 350,000 units

b. between 350,001 and 400,000 units

c. between 400,001 and 450,000 units

d. between 450,001 and 500,000 units
Answer #1

Fixed cost = 1000000 + 3000000 = 4000000 = FC

Variable cost = 3+1 = 4$ = VC

Selling price = 13$ = SP

Break even point = ( FC )/ (SP - VC)

Break even point = (4000000)/(13-4)

BEP = (4000000/9) = 444444.44 unit

Break even point is in between = 400001 to 450000 units

Answer: C